OpenStudy (anonymous):

Tino conducted a survey of all the eighth graders in his school for his math project.He asked each student how much he or she spent on school supplies at the beginning of the year (rounded to the nearest dollar) he also caculated the average amount spent and found it to be $15.50 unfortunately,it rained on the day he brought the graph to turn it in,and the last bar got smudged.What is the height of the smudged are?

OpenStudy (anonymous):

go to this link and and scroll down youll see the graph

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

take a weighted average

OpenStudy (amistre64):

or ... sum xP(x) is the same thing

OpenStudy (anonymous):

did you find the problem on the link @amistre64

OpenStudy (amistre64):

\[\frac{a\alpha+b\beta+c\gamma+d\delta+...+k\kappa}{a+b+c+d+...+k}\]

OpenStudy (amistre64):

i found it, we need to take a weighted average

OpenStudy (anonymous):

um im in 8th grade

OpenStudy (anonymous):

can i tell you what my teacher gave me

OpenStudy (anonymous):

for an equation @amistre64

OpenStudy (amistre64):

only if its sensible ...

OpenStudy (anonymous):

total spent + 20x -------------- 18(12)=216 20(13)=260 number of kids + x

OpenStudy (amistre64):

ideally, if we add up all the amounts that each kid gave, and divide by how many kids gave an answer .... we have an average. 18 kids gave 12: 12+12+12+12+12+ ... or 18(12) 20 kids gave 13: 13+13+13+ .. or simply 13(20) 21 kids gave 14: 14(21) see how its working?

OpenStudy (anonymous):

yes thats what he wanted us to do but i dont see how its working @amistre64

OpenStudy (amistre64):

\[\frac{18(13)+ 20(14)+ 21(15)+ 22(16)+ 18(17)+ 16(18)+ 8(19)+ 10(20)+ x(21)}{18+20+21+22+18+16+8+10+x}\\ =15.50\]

OpenStudy (amistre64):

i mis saw something in there trying to flip between pages

OpenStudy (amistre64):

but thats the basic premise.

OpenStudy (amistre64):

oh, we start at 12, not 13

OpenStudy (amistre64):

\[\frac{18(12)+ 20(13)+ 21(14)+ 22(15)+ 18(16)+ 16(17)+ 8(18)+ 10(19)+ x(20)}{18+20+21+22+18+16+8+10+x}\] simplify that for me

OpenStudy (anonymous):

this might take awhile @amistre64

OpenStudy (anonymous):

hold on

OpenStudy (amistre64):

while your pondering it ... do you agree that an average is simply the sum of the number of items, divided by how many items there are?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then thats really all we are doing here. but being smarter about it 18 kids gave $12 if we add 12 to itself 18 times ... thats the definition of multiplication: 18(12) the top part is just summing up all the money that was given by taking the frequency of each dollar amount instead of adding them up one by one

OpenStudy (amistre64):

the bottom is therefore how many kids gave in the first place ... 18 gave this 20 gave that 21 gave those 22 gave something else .... the sum of the number of kids is the bottom

OpenStudy (anonymous):

2024x ------- = 15. 133z

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

1994 + 20x ----------- = 15.50 133 + x solve for x

OpenStudy (anonymous):

how did you get 1994

OpenStudy (anonymous):

i got 2024

OpenStudy (anonymous):

did i add wrong

OpenStudy (anonymous):

oh k

OpenStudy (amistre64):

there was not way i was going to chk that by hand :)

OpenStudy (anonymous):

i put it in the caculator and got 2014

OpenStudy (amistre64):

you might need new batteries

OpenStudy (amistre64):

so lets assume we did this right: 18(12)+ 20(13)+ 21(14)+ 22(15)+ 18(16)+ 16(17)+ 8(18)+ 10(19) + 20x 1994 + 20x on top and 18+20+21+22+18+16+8+10 + x 133 + x on the bottom

OpenStudy (anonymous):

its 2,014 i used the google caculator but i didnt add the ask just 20

OpenStudy (anonymous):

but its 2014x

OpenStudy (amistre64):

using m for the mean 1994 + 20x = m(133+x) 1994 + 20x = 133m + mx 20x - mx = 133m - 1994 x(20 - m) = 133m - 1994 133m - 1994 x = ------------ 20 - m

OpenStudy (amistre64):

you can add unlike terms ... 20x is not able to be added for starters. it has to be worked in seperately

OpenStudy (amistre64):

if you want to use your findings, feel free ....

OpenStudy (anonymous):

1994 + 20x ----------- = 15.50 133 + x

OpenStudy (amistre64):

yep .... i used m for a simpler run thru

OpenStudy (anonymous):

im confused on wht to do

OpenStudy (amistre64):

using m for the mean 1994 + 20x = m(133+x) 1994 + 20x = 133m + mx 20x - mx = 133m - 1994 x(20 - m) = 133m - 1994 133m - 1994 x = ------------ 20 - m

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

you wnat me to solve

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

yes, i want you to solve

OpenStudy (anonymous):

do i have to use the m

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what does m equal in this case? i defined it to make my life simpler

OpenStudy (anonymous):

mean i was just asking i thought i could use x

OpenStudy (amistre64):

x is what is not known :) m is known, just easier for me to work with when moving things about

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

15.0

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

x = 15 is what i get as well

OpenStudy (anonymous):

so that was it

OpenStudy (amistre64):

with any luck yes :) i would not have attempted that by hand, which is why i trust the wolf on it

OpenStudy (anonymous):

oh thank you finally finished

OpenStudy (anonymous):

thank you so much @amistre64

OpenStudy (amistre64):

youre welcome :)